How do I go from knowing the rational roots (2, -3, -6) to the equation

Answer by richwmiller(17219)

set up factors
(x-2)*(x+3)*(x+6)=0
notice that when each factor is set to 0 they work out to the roots
x-2=0
x=2
etc
now multiply out the factors
(x-2)*(x+3)*(x+6)
use foil to multiply (x+3)*(x+6)
x^2+9 x+18
now multiply each term by x and then do the same for -2
combine and add them together to get
x^3+7x^2-36 = 0